(rs)radical sign


(rs)180
My answer is 13.42 is this correct?

Found 2 solutions by chitra, praseena:
Answer by chitra(359)

Solution:
Yes, you answer is prefectly correct!!
The given question is:
rs(180) which is equal to 13.416407
So rounding off to two digits after the decimal point, we get:
rs(180) = 13.42
Regards,
Chitra

Answer by praseena(37)

your answer is correct.
here is the solution;
Find the two perfect square numbers in which 180 lies between.
here we have the perfect numbers
13^2=169 and 14^2=196
so square root of 180 lies between 13 and 14

Now Divide 180 by 13. (you can also divide 180 by 14 and take the average of
14 and that answer and then you can follow can same method)


180/13 = 13.84(you can round off your answer)
Now find Average 0f 13.84 and 13.

(13.84 + 13)/2 = 26.84/2 =13.42

Note: There are a number of ways to calculate square roots without a calculator. This is only one of them.